A spring stretches 30 cm to hold up a 4 kg mass. How far would the spring stretch if it is used to accelerate a 9 kg mass upward at 5 m/s2.
1
Expert's answer
2020-11-25T10:50:44-0500

The force required to accelerate 9 kilos to 5 m/s2 is


"F=ma."

According to Hooke's law, the spring needs to be stretched for


"x=\frac Fk=\frac{ma}{k}."

The stiffness coefficient can be found from the experiment with 30 cm compression and 4 kg mass:


"k=\frac{F_e}{x_e}=\frac{m_eg}{x_e}.\\\space\\
x=x_e\frac{ma}{m_eg}=0.3\frac{9\cdot5}{4\cdot9.8}=0.34\text{ m}."